OpenStudy (anonymous):

Simplify the expression: [(18 – 6) • 3 + 1] • 7

OpenStudy (anonymous):

[(12)*3+1]*7= [36+1]*7= 37*7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yiur welcome